OpenStudy (anonymous):

If cos (2m)=.11 what is (cos m)^2

OpenStudy (anonymous):

Use the cosine half-angle identity:\[\cos ^{2}m = \frac{ 1 + \cos 2m }{ 2 }\]

OpenStudy (anonymous):

2.22 is that correct

OpenStudy (anonymous):

I'm not sure how you got that. It's not correct, but if you show me your work, step by step, I'm sure I can see where you went wrong.

OpenStudy (anonymous):

1.11

OpenStudy (anonymous):

I was trying to add wrong number

OpenStudy (anonymous):

1+2(.11)/2 =

OpenStudy (anonymous):

Okay, I see where you went wrong. cos 2m is equal to .11 , not cos m = .11\[\frac{ 1 + .11 }{ 2 }\]

OpenStudy (anonymous):

Also,\[\cos ^{2}m = (\cos m)^{2}\]

OpenStudy (anonymous):

So, it comes down to 0.555 for (cos m)^2

OpenStudy (anonymous):

.11/2

OpenStudy (anonymous):

No, 1.11/2

OpenStudy (anonymous):

sorry thank you so much

OpenStudy (anonymous):

np. You did well.

OpenStudy (anonymous):

your really smart I guess you already know that Huh

OpenStudy (anonymous):

My dog is about the only one who thinks I'm smart, and that's only because I know where the food is. I secretly think he knows where the food is, too! :-)

OpenStudy (anonymous):

Lol should put up a picture of sheldon from big bang theory

OpenStudy (anonymous):

Sheldon rocks! Kaboom! (elevator episode)

OpenStudy (anonymous):

So do you lifesaver ;D

OpenStudy (anonymous):

Have a good one!